ONLY choose the mcq don't explain solve the exam in 30 mins Four wires are joined...

  1. ONLY choose the mcq don't explain solve the exam in 30 mins

  2. Four wires are joined together at a point P. Current 15 A flows towards P from first wire. A current of 10 A flows away from P through the second wire. Through the third wire, a current of 8 A flows towards P. How much current flows through the fourth wire and in which direction?

    A.

    17 A away from P

    B.

    17 A towards P

    C.

    13 A away from P

    D.

    13 A towards P

  3. Choose the correct statement regarding resistance combination.

    A.

    When resistors are connected in parallel voltage remains the same across each resistor and the total resistance decreases

    B.

    When resistors are connected in parallel current remains the same across each resistor and the total resistance increases

    C.

    When resistors are connected in parallel voltage remains the same across each resistor and the total resistance increases

    D.

    When resistors are connected in parallel current remains the same across each resistor and the total resistance decreases

  4. 5 light bulbs of resistance 200 ohms are connected in parallel. The effective resistance of the combination is

    A.

    1000 ohms

    B.

    200 ohms

    C.

    500 ohms

    D.

    40 ohms

  5. Unit of k the constant in Coulomb's Law is

    A.

    Nm2C-2

    B.

    N-1m2C-2

    C.

    N2mC-2

    D.

    N-1m-2C2
Answer #1

1.

Answer is

C.

13 A away from P

From Kirchoff's Junction rule

I1+I2+I3+I4=0

-15+10-8+I4=0

I4=13 A (away from P)

2.

Answer is

When resistors are connected in parallel voltage remains the same across each resistor and the total resistance decreases

3.

Answer is

D.

40 ohms

equivalent resistance

R =(1/200 +1/200 +1/200 +1/200 +1/200)-1

R=200/5 =40 ohms

4,

Answer is

A.

Nm2C-2

Since Coloumb Constant is given

K=8.99*109 Nm2C-2